javonarnold:

Victor wrote the following statement: "You can only draw one unique isosceles triangle that contains an angle of 85°." Which statement is true? (1 point) Victor is correct, because only one unique triangle can be drawn with the given information. Victor is incorrect, because more than one triangle can be drawn with the given information. Victor is incorrect, because the triangle described cannot be drawn with the given information. None of the above.

javonarnold:

A circle is cut from a square piece of cloth, as shown: A square, one side labeled as 34 inches, has a circle inside it. The circle touches all the sides of the square. The portion of the square outside the circle is shaded. How many square inches of cloth are cut from the square? (π = 3.14) (1 point) 907.46 in2 986.32 in2 1,024.00 in2 1,156.00 in2

jhonyy9:

so like a first step you need to know what is the propertie of an isoscele triangle ? how are the lateral sides and the angles corresponding to these sides ? @AZ opinion pls. ? ty.

surjithayer:

85+85=170+10=180 180-85=95 95/2=47.5 so two isosceles triangles can be drawn. \[85 ^\circ,85^\circ,10^\circ\] \[85^\circ,47.5^\circ,47.5^\circ\] It is B.